Question 1: (20 pts) A university wants to find out whether an Englsh gmogram heligs fheir udets m skills. They design a if the students soore is hi test to checic for te a) Can the Final Exam
Can the univ English i Student Before English Program Afer 16 21 19 16 24 20 17 29 17 - 20 20 12 13 14 15 14 19 16 18 26
Answer #1
Before After d = After - Before
18 22 4
16 17 1
21 25 4
19 16 -3
22 24 2
17 20 3
24 29 5
17 20 3
18 20 2
23 19 -4
16 15 -1
14 15 1
19 26 7
16 18 2
18 18 0
d-bar = 1.733333333
s = 2.914659173

Data:     

n = n1 = n2 = 15    

d-bar = 1.733333333    

s (of d) = 2.914659173    

Hypotheses:     

Ho: d-bar ≤ 0    

Ha: d-bar > 0    

Decision Rule:     

α = 0.05    

Degrees of freedom = 15 - 1 = 14

Critical t- score = 1.761310115   

Reject Ho if t > 1.761310115    

Test Statistic:     

SE = s/√n = 2.91465917308321/√15 = 0.752561762   

t = d-bar/SE = 1.73333333333333/0.752561762481463 = 2.303243959   

p- value = 0.0185606    

Decision (in terms of the hypotheses):    

Since 2.303243959 > 1.761 we reject Ho and accept Ha

Conclusion (in terms of the problem):    

There is sufficient evidence to conclude that the program is helpful.